OpenStudy (anonymous):

simplify (x^1/4)^4/9

jimthompson5910 (jim_thompson5910):

use the rule (a^b)^c = a^(b*c)

OpenStudy (anonymous):

x^7/36?

jimthompson5910 (jim_thompson5910):

how are you getting 7?

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

is 1/4 really 7/4?

OpenStudy (anonymous):

over 13

jimthompson5910 (jim_thompson5910):

all you need to do is multiply the exponents 1/4 and 4/9

jimthompson5910 (jim_thompson5910):

make sure to reduce as much as possible

OpenStudy (anonymous):

would it be 4/13?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

25/36

jimthompson5910 (jim_thompson5910):

hopefully you're not just guessing the answer choices

OpenStudy (anonymous):

9*4=36

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

the numerators multiply to ?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so it's 4/36 which reduces to what?

OpenStudy (anonymous):

1/9

jimthompson5910 (jim_thompson5910):

thats the exponent of the answer

OpenStudy (anonymous):

how abou ths one? Simplify (7 – 2i) – (3 + 9i)

jimthompson5910 (jim_thompson5910):

Hint: (7 – 2i) – (3 + 9i) = 7 – 2i – 3 - 9i

OpenStudy (anonymous):

4-11i

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

don't forget, when you're distributing that negative through, it has to go to each term inside

OpenStudy (anonymous):

solve sqrt x+3-3=x

jimthompson5910 (jim_thompson5910):

sqrt(x+3) - 3 = x sqrt(x+3) = x+3 x+3 = (x+3)^2 x+3 = x^2+6x+9 I'll let you finish

OpenStudy (anonymous):

x=-2 x=-3

jimthompson5910 (jim_thompson5910):

now check those potential answers

OpenStudy (anonymous):

-3 doesnt work

jimthompson5910 (jim_thompson5910):

whys that

OpenStudy (anonymous):

wait -3=-3

jimthompson5910 (jim_thompson5910):

so that confirms -3

jimthompson5910 (jim_thompson5910):

what about -2

OpenStudy (anonymous):

the both work

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

sometimes only one works or sometimes none work, so that's why you need to check both

OpenStudy (anonymous):

Explain, in complete sentences, how you might show that radicals should be simplified before being added or subtracted to another radical.

jimthompson5910 (jim_thompson5910):

what do you think on this one

OpenStudy (anonymous):

i honestly dont knoe

jimthompson5910 (jim_thompson5910):

well think about it for a sec

jimthompson5910 (jim_thompson5910):

how would you simplify any radical

OpenStudy (anonymous):

square root

jimthompson5910 (jim_thompson5910):

yes how would you simplify those

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

i dont understand

jimthompson5910 (jim_thompson5910):

how would you simplify something like sqrt(48)

OpenStudy (anonymous):

square 48

jimthompson5910 (jim_thompson5910):

If you're not sure, this page explains how to simplify radicals http://www.regentsprep.org/Regents/math/ALGEBRA/AO1/Lsimplify.htm